40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 22 19:53:01 2013, Jenne, Update, Locking, PR-flat cavity status - not locked 
    Reply  Wed Jan 23 19:05:16 2013, Jenne, Update, Locking, PR-flat cavity status - locks! 
       Reply  Wed Jan 23 20:46:46 2013, Zen Master, Update, Locking, PR-flat cavity status - locks! yinYang.png
Message ID: 7928     Entry time: Tue Jan 22 19:53:01 2013     Reply to this: 7931
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PR-flat cavity status - not locked 

The PR-flat cavity is flashing, although not locked.  I am too hungry to continue right now.

I put the FI_Back camera on a tripod, looking at the back of the Faraday.  The beam that Jamie and I were working with on Friday was clipped going back through the Faraday.  I twiddled the TT2 and PRM pointing such that the beam is retroreflecting, and getting back through the Faraday, and the cavity is still flashing.  I then redid the REFL path on the AS table a little bit.  The beam is currently going to the REFL camera, as well as REFL11 and REFL55. 

Some notes about the AS table:  The Y1 separating the main REFL beam from the REFL camera beam was mounted 90 degrees (rotated about the beam's axis) from what it should be.  I fixed it, so that the straight-through beam that goes to the camera is not clipped by the edge of the mount.  The reason (I think) this mirror was mounted backwards is that when mounted correctly, the back of the mount and the knobs interfere with the AS beam path.  I solved this by rotating the first out-of-vac REFL mirror a small amount so that the REFL and AS beams are slightly more separated. 

I am not seeing any nice PDH signal on dataviewer, so I went to check the signal path for the PDs.  The 11MHz marconi is on and providing RF, the EOM is plugged in to 11, 55 and 29.5 signals (no aux cavity scan cables are plugged in).  Both of the RF Alberto boxes are on.  I measured the RF output of both REFL11 and REFL55, although after the fact I realized that I was BAD, and had not found a 'scope that lets me change the input impedance to 50 ohms.  BAD grad student.  However, since I have numbers, I will post them, despite their being not quite correct:

284mVpp at 11MHz out of REFL11.  This is -6.9dBm

2mVpp at 55MHz out of REFL55, measured by 'scope

So, I can clearly see the 11MHz on the 'scope, and can see a very noisy, small 55MHz signal on the 'scope.  I need to think over dinner about what level of signal we should be sending to the demod boards, and whether or not I need more power coming out of the RFPDs.  There is a wave plate and PBS before beam goes to any of the REFL PDs, presumably to ensure that none of them get fried when we're at high power.  If I need more signal, I suspect I can rotate the wave plate and let more light go to the diodes.

ELOG V3.1.3-